# Deep Learning-based Intrusion Detection Systems: A Survey

### v2

@@ Abstract.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 4ccs: Security and privacy Intrusion detection systemsccs: Computing methodologies Machine learningccs: General and reference Surveys and overviews
@@ 2. Survey Methodology
To start our literature review, we selected several popular literature databases, including Web of Science (Analytics, 1997), IEEE Xplore (IEEE, 2000), and Scopus (Elsevier, 2021), as the search engine. For search keywords, we determined from generalized terms associated with DL-IDS, such as intrusion detection system, attack investigation, anomaly detection, threat detection, Advanced Persistent Threats, data provenance analysis, forensic analysis, causality analysis, log collection, log compression, log parsing, log storage, and log summarization. Then, we employed Connected Papers (Papers, 2020), a visual tool that assists researchers in finding relevant academic papers, to ensure that we did not overlook the typical related literature. Since the found literature is numerous and rather generalized for the DL-IDS scope, we carefully checked their topics and prioritized only academic papers that are related to DL. Finally, all these papers were filtered based on the impact factors of their published journals or academic conferences, leaving us a total of 113 papers.
We identified a few venues that have published many significant papers in the field of DL-IDS, such as Security, S&P, CCS, NDSS, TIFS, TDSC, ICSE, ASE, ESEC/FSE, TSE, EuroSys, NSDI, OSDI, SOSP, ATC, ICDE, TKDE, SIGMOD, KDD, WWW, and IJCAI. We broadly divide them into four categories: security, software, system, and data. The distribution of these papers with their published years is reported in Figure 3.
@@ 4.1. Data Collection
The first step of DL-IDS is to collect useful data from raw data. Raw data indicates records that document events, activities, and operations that occur within a system, application, or network (a.k.a., logs), represented by audit logs or application logs within hosts, or network traffic between hosts. By collecting useful logs, DL-IDS is allowed to monitor the health condition and operational status of information systems (Zhang et al., 2023b). Common attributes of logs include timestamp, event type, subject, object, description, etc.
@@ 4.2.2. Log Compression Algorithms
Different from natural language data, log data usually has specific structures and formal expressions that help further compression. Yao et al. (Yao et al., 2021) propose LogBlock, which obtains small log blocks before compression and then uses a generic compressor to compress logs. Liu et al. (Liu et al., 2019) propose Logzip, which employs clustering algorithms to iteratively extract templates from raw logs and then obtain coherent intermediate representations for compressing logs. Rodrigues et al. (Rodrigues et al., 2021) propose the lossless compression tool CLP, aiming to quickly retrieve log data while meeting compression requirements. CLP proposes to combine domain-specific compression and search with a generic lightweight compression algorithm. Li et al. (Li et al., 2024d) conduct empirical research on log data and propose LogShrink to overcome their observed limitations by leveraging the commonality and variability of log data. LogBlock (Yao et al., 2021) is designed to help existing jobs perform better. It reduces duplicate logs by preprocessing log headers and rearranging log contents, thereby improving the compression ratio of log files. LogReducer (Yu et al., 2023) is a framework that combines log hotspot identification and online dynamic log filtering. Its non-intrusive design significantly reduces log storage and runtime overhead. $\mu$Slope (Wang et al., 2024) is a compression and search method for semi-structured log data. It achieves efficient storage and query performance through data segmentation, pattern extraction, and index-free design. Denum (Yu et al., 2024) significantly improves log compression rates by optimizing the compression of digital tokens in logs. It is an efficient log compression tool suitable for scenarios where you need to save storage space or transmission bandwidth.
@@ 5.1.1. Offline Graph Summarization
Offline graph summarization requires historical log data to provide global knowledge, which extracts log data from persistent storage, summarizes the data, and pushes back the summarized data to the persistent storage. In 2011, Xie et al. (Xie et al., 2011) take inspiration from web graphs to summarize provenance graphs. They argue that provenance graphs have similar organizational structure and characteristics to web graphs, such as locality, similarity, and consecutiveness. BEEP (Lee et al., 2013a) is developed based on the fact that a long-running execution can be partitioned into individual units. BEEP reverse engineers application binaries and instructions to perform selective logging for unit boundaries and unit dependencies. LogGC (Lee et al., 2013b) is a summarized audit log system that can be invoked at any time during the system execution. Xu et al. (Xu et al., 2016) propose an aggregation algorithm PCR that preserves event dependencies during log data reduction. They further propose an algorithm named PCAR that utilizes domain knowledge to conduct graph summarization. Hossain et al. (Hossain et al., 2018) propose two dependency-preserving graph summarization approaches, FD and SD. FD is allowed to keep backward and forward forensic analysis results. SD preserves the results of common forensic analysis, which runs backward to find the entry points of intrusions and then runs forward from these points to unveil their impacts. LogApprox (Michael et al., 2020) aims to summarize the most space-intensive events found in logs, namely file I/O activity, which can account for up to 90% of the log content.
@@ 5.2. Attack Detection
Log Graph threaTrace (Wang et al., 2022) 2022 GraphSAGE Online Node
Neural ProGrapher (Yang et al., 2023) 2023 graph2vec Online Subgraph
Network MAGIC (Jia et al., 2024) 2024 GAT Online Node, Subgraph
Flash (Rehman et al., 2024) 2024 GraphSAGE Online Node
@@
Application Sequence LogAnomaly (Meng et al., 2019) 2019 template2vec, LSTM Online Log Entry
Log Neural LogC (Yin et al., 2020) 2020 LSTM Online Log Entry
Network NeuralLog (Le and Zhang, 2021) 2021 BERT - Log Entry
PLELog (Yang et al., 2021) 2021 Attention GRU Online Log Entry
@@
Traffic Whisper (Fu et al., 2021) 2021 K-Means - Host
Traditional SigML++ (Trivedi et al., 2023) 2023 ANN - Encrypted Log
Learning OADSD (Zhang et al., 2023a) 2023 Isolation Forest Online Packet
@@
Graph and ERNN (Zhao et al., 2022b) 2022 ERNN - Flow
Sequence Euler (King and Huang, 2023) 2023 GNN, RNN - Flow
Neural pVoxel (Fu et al., 2023) 2023 - - Packet, Flow
Network NetVigil (Hsieh et al., 2024) 2024 E-GraphSage - Flow
Exosphere (Fu et al., 2024) 2024 CNN - Packet
DFNet (Zhao et al., 2024a) 2024 DFNet - Packet
@@ 5.2.1. Audit Log-based Detectors
Traditional learning-based detectors refer to those that utilize naive machine learning techniques. StreamSpot (Manzoor et al., 2016) is a clustering-based anomaly detection that tackles challenges in heterogeneity and streaming nature. Unicorn (Han et al., 2020) is a real-time intrusion detector that efficiently constructs a streaming histogram to represent the history of system executions. The counting results within the histogram are updated immediately if new edges (or events) occur. DistDet (Dong et al., 2023) is a distributed detection system that builds host models in the client side, filters false alarms based on their semantics, and derives global models to complement the host models.
@@
GNN is demonstrated to do well in processing provenance graphs (Yang et al., 2023; Jia et al., 2024; Rehman et al., 2024; Zengy et al., 2022; Wang et al., 2022). ProGrapher (Yang et al., 2023) extracts temporal-ordered provenance graph snapshots from the ingested logs, and applies whole graph embedding and sequence-based learning to capture rich structural properties of them. The key GNN technique leveraged by ProGrapher is graph2vec. ShadeWatcher (Zengy et al., 2022) is a recommendation-guided intrusion detector using provenance graphs. It borrows the recommendation concepts of user-item interactions into security concepts of system entity interactions and analyzes cyber threats in an automated and adaptive manner. threaTrace (Wang et al., 2022) emerges as an online approach dedicated to detecting host-based threats at the node level. Its GNN model is a tailored GraphSAGE (Hamilton et al., 2017) for learning rich contextual information in provenance graphs. MAGIC (Jia et al., 2024) leverages Graph Attention Network (GAT) (Velickovic et al., 2017) as its graph representation module. MAGIC employs masked graph representation learning to incorporate the capability of pre-training. It can adapt to concept drift with minimal computational overhead, making it applicable to real-world online APT detection. Flash (Rehman et al., 2024) is a comprehensive and scalable approach on data provenance graphs to overcome the limitations in accuracy, practicality, and scalability. Flash incorporates a novel adaptation of a GNN-based contextual encoder to encode both local and global graph structures into node embeddings efficiently. R-caid (Goyal et al., 2024) first incorporates root cause analysis into PIDS. Before training GNNs, R-caid links nodes to their root causes to build a new graph, intending to prevent it from mimicry and evasion attacks. Argus (Xu et al., 2024b) finds the performance of the prior IDS is questionable on large scale. It thus devises a form of discrete temporal graph and uses encoder-decoder unsupervised learning to detect different types of attacks.
@@ 5.2.2. Application Log-based Detectors
For traditional learning, Wei et al. (Xu et al., 2009) propose a general methodology to mine rich semantic information in console logs to detect large-scale system problems. Bodik et al. (Bodik et al., 2010) leverage a logistic regression model on a new and efficient representation of a datacenter’s state called fingerprint to detect previously seen performance crises in that datacenter.
@@
Due to the similarity between application logs and natural language texts, sequence neural networks such as Recurrent Neural Network (Hochreiter and Schmidhuber, 1997) and Transformer (Vaswani et al., 2017; Devlin et al., 2019) are widely employed. DeepLog (Du et al., 2017) employs LSTM to model system logs as natural language sequences. It is able to automatically learn benign log patterns and detect anomalies when there is a deviation between log patterns and the trained model. LogRobust (Zhang et al., 2019) finds previous methods do not work well under the close-world assumption and utilizes an attention-based LSTM model to handle unstable log events and sequences. LogAnomaly (Meng et al., 2019) identifies previous studies tend to cause false alarms by using indexes rather than semantics of log templates. Empowered by a novel, simple yet effective method termed template2vec, LogAnomaly is proven to successfully detect both sequential and quantitive log anomalies simultaneously. LogC (Yin et al., 2020) is a new log-based anomaly detection approach with component-aware analysis. It feeds both log template sequences and component sequences to train a combined LSTM model for detecting anomalous logs. NeuralLog (Le and Zhang, 2021) targets the performance caused by log parsing errors such as out-of-vocabulary words and semantic misunderstandings and employ BERT to perform neural representation. PLELog (Yang et al., 2021) is a semi-supervised anomaly detection approach that can get rid of time-consuming manual labeling and incorporate the knowledge on historical anomalies. SpikeLog (Qi et al., 2023) adopts a weakly supervised approach to train an anomaly score model, with the objective of handling a more reasonable premise scenario where a large number of logs are unlabeled. LogCraft (Zhang et al., 2024) is an end-to-end unsupervised log anomaly detection framework based on automated machine learning, which mitigates the cost of understanding datasets and makes multiple attempts for building algorithms.
@@ 5.2.3. Network Traffic-based Detectors
Given the fact that network traffic is usually encrypted for secure communications, feature engineering-guided machine learning is widely applied in NIDS. Whisper (Fu et al., 2021) pays attention to both high accuracy and high throughput by utilizing frequency domain features. SigML++ (Trivedi et al., 2023) is an extension of SigML for supervised anomaly detection approach. SigML++ employs Fully Homomorphic Encryption and Artificial Neural Network (ANN) for detection, resulting in execution without decrypting the logs. OADSD (Zhang et al., 2023a) achieves task independently and has the ability of adapting to the environment over SD-WAN by using On-demand Evolving Isolation Forest. LtRFT (Tang et al., 2023) innovatively introduces Learning-To-Rank scheme for mitigating the low-rate DDoS attacks targeted at flow tables.
@@
In network traffic, packets consist of various contents and their flows can be represented as graphs. As a result, both graph neural network and sequence neural network are adopted in NIDS. Kitsune (Mirsky et al., 2018) is a plug and play NIDS that is allowed to detect attacks efficiently on the local network without supervision. It alleviates the problem that network gateways and router devices simply do not have the memory or processing power. MT-FlowFormer (Zhao et al., 2022a) is a semi-supervised framework to mitigate the lack of a mechanism for modeling correlations between flows and the requirement of a large volume of manually labeled data. I2RNN (Song et al., 2022) is an incremental and interpretable RNN for encrypted traffic classification, which can be efficiently adapted for incremental traffic types. ERNN (Zhao et al., 2022b) represents error-resilient RNN, which is a robust and end-to-end RNN model specially designed against network-induced phenomena. Euler (King and Huang, 2023) accelerates the most memory-intensive part, message-passing stage within GNN, with several concurrently-executed replicated GNNs. pVoxel (Fu et al., 2023) is an unsupervised method that proposes to leverage point cloud analysis to reduce false positives for the previous NIDS such as Whisper and Kitsune without requiring any prior knowledge on the alarms. NetVigil (Hsieh et al., 2024) is specially designed for east-west traffic within data center networks. It utilizes E-GraphSage and contrastive learning techniques to strengthen its resilience. Exosphere (Fu et al., 2024) detects flooding attacks by analyzing packet length patterns, without investigating any information in encrypted packets. DFNet (Zhao et al., 2024a) is a DDoS prevention paradigm denoted by preference-driven and in-network enforced shaping. RFH-HELAD (Zhong et al., 2024) consists of a $K$ classification model based on a deep neural network and a $K+1$ classification combining GAN and Deep kNN for detecting anomalies in network traffic.

### v3

@@ Abstract.
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 1journalnumber: 1article: 1publicationmonth: 10ccs: Security and privacy Intrusion detection systemsccs: Computing methodologies Machine learningccs: General and reference Surveys and overviews
@@ 2. Survey Methodology
To start our literature review, we selected several popular literature databases, including Web of Science (Analytics, 1997), IEEE Xplore (IEEE, 2000), and Scopus (Elsevier, 2021), as the search engine. For search keywords, we determined from generalized terms associated with DL-IDS, such as intrusion detection system, attack investigation, anomaly detection, threat detection, Advanced Persistent Threats, data provenance analysis, forensic analysis, causality analysis, log collection, log compression, log parsing, log storage, and log summarization. Then, we employed Connected Papers (Papers, 2020), a visual tool that assists researchers in finding relevant academic papers, to ensure that we did not overlook the typical related literature. Since the found literature is numerous and rather generalized for the DL-IDS scope, we carefully checked their topics and prioritized only academic papers that are highly related. Finally, all these papers were filtered based on the impact factors of their published journals or academic conferences, leaving us a total of 131 papers.
We identified a few venues that have published many significant papers in the field of DL-IDS, such as Usenix Security, S&P, CCS, NDSS, TIFS, TDSC, ICSE, ASE, ESEC/FSE, TSE, OSDI, NSDI, EuroSys, SOSP, ATC, ICML, KDD, WWW, TKDE, ICDE, and SCIS. We broadly divide them into five categories: security, software, system, data, and interdiscipline. The distribution of these papers with their published years is reported in Figure 3.
@@ 4.1. Data Collection
The first step of DL-IDS is to collect useful data from raw data. Raw data indicates records that document events, activities, and operations that occur within a system, application, or network (a.k.a., logs), represented by audit logs or application logs within hosts, or network traffic between hosts. By collecting useful logs, DL-IDS is allowed to monitor the health condition and operational status of information systems (Zhang et al., 2023b; Ma et al., 2018). Common attributes of logs include timestamp, event type, subject, object, description, etc.
@@ 4.2.2. Log Compression Algorithms
[…]
@@ 5.1.1. Offline Graph Summarization
Offline graph summarization requires historical log data to provide global knowledge, which extracts log data from persistent storage, summarizes the data, and pushes back the summarized data to the persistent storage. In 2011, Xie et al. (Xie et al., 2011) take inspiration from web graphs to summarize provenance graphs. They argue that provenance graphs have similar organizational structure and characteristics to web graphs, such as locality, similarity, and consecutiveness. BEEP (Lee et al., 2013a) is developed based on the fact that a long-running execution can be partitioned into individual units. BEEP reverse engineers application binaries and instructions to perform selective logging for unit boundaries and unit dependencies. LogGC (Lee et al., 2013b) is a summarized audit log system that can be invoked at any time during the system execution. Xu et al. (Xu et al., 2016) propose an aggregation algorithm PCR that preserves event dependencies during log data reduction. They further propose an algorithm named PCAR that utilizes domain knowledge to conduct graph summarization. Hossain et al. (Hossain et al., 2018) propose two dependency-preserving graph summarization approaches, FD and SD. FD is allowed to keep backward and forward forensic analysis results. SD preserves the results of common forensic analysis, which runs backward to find the entry points of intrusions and then runs forward from these points to unveil their impacts. LogApprox (Michael et al., 2020) aims to summarize the most space-intensive events found in logs, namely file I/O activity, which can account for up to 90% of the log content. TeRed (Li et al., 2025) employs unit tests to learn the system’s normal behavior patterns for reducing provenance graphs, allowing it not to impact attack detection and investigation.
@@ 5.2. Attack Detection
Log threaTrace (Wang et al., 2022) 2022 GraphSAGE Online Node
Graph ProGrapher (Yang et al., 2023) 2023 graph2vec Online Subgraph
Neural MAGIC (Jia et al., 2024) 2024 GAT Online Node, Subgraph
Network Flash (Rehman et al., 2024) 2024 GraphSAGE Online Node
@@
Application LogAnomaly (Meng et al., 2019) 2019 template2vec, LSTM Online Log Entry
Log Sequence LogC (Yin et al., 2020) 2020 LSTM Online Log Entry
Neural NeuralLog (Le and Zhang, 2021) 2021 BERT - Log Entry
Network PLELog (Yang et al., 2021) 2021 Attention GRU Online Log Entry
@@
Tweezers(Cui et al., 2024) 2024 GATv2, BERTweet Online Log Entry
LogSer (Chai et al., 2024) 2024 BERT Online Log Entry
LogDLR(Zhou et al., 2025) 2025 Transformer, SBERT Online Log Entry
Traffic NetPro (Li et al., 2017) 2017 Merkle Hash Tree Online Route
Log CATH (Guo et al., 2019) 2019 Cusp Model Online Flow
Traditional Whisper (Fu et al., 2021) 2021 K-Means - Host
Learning SigML++ (Trivedi et al., 2023) 2023 ANN - Encrypted Log
OADSD (Zhang et al., 2023a) 2023 Isolation Forest Online Packet
@@
ERNN (Zhao et al., 2022b) 2022 ERNN - Flow
Euler (King and Huang, 2023) 2023 GNN, RNN - Flow
Graph and pVoxel (Fu et al., 2023) 2023 - - Packet, Flow
Sequence NetVigil (Hsieh et al., 2024) 2024 E-GraphSage - Flow
Neural Exosphere (Fu et al., 2024) 2024 CNN - Packet
Network DFNet (Zhao et al., 2024b) 2024 DFNet - Packet
@@
ReTrial (Zhao et al., 2024a) 2024 Bayesian Inference Online Flow
HEN (Wei et al., 2024) 2024 AE-LSTM - Packet, Flow
TCG-IDS (Wu et al., 2025b) 2025 TGN Online Flow
A-NIDS(Zha et al., 2025) 2025 Stacked CTGAN Online Flow
GTAE-IDS(Ghadermazi et al., 2025) 2025 Graph Transformer Online Packet, Flow
@@ 5.2.1. Audit Log-based Detectors
Traditional learning-based detectors refer to those that utilize naive machine learning techniques. StreamSpot (Manzoor et al., 2016) is a clustering-based anomaly detection that tackles challenges in heterogeneity and streaming nature. Unicorn (Han et al., 2020) is a real-time intrusion detector that efficiently constructs a streaming histogram to represent the history of system executions. The counting results within the histogram are updated immediately if new edges (or events) occur. DistDet (Dong et al., 2023) is a distributed detection system that builds host models in the client side, filters false alarms based on their semantics, and derives global models to complement the host models. Velox (Bilot et al., 2025) derives from Orthrus and replaces the complex TGN-based encoder with a simple fully-connected network (FCN), leading to a lightweight and efficient neural network.
@@
GNN is demonstrated to do well in processing provenance graphs (Yang et al., 2023; Jia et al., 2024; Rehman et al., 2024; Zengy et al., 2022; Wang et al., 2022). ProGrapher (Yang et al., 2023) extracts temporal-ordered provenance graph snapshots from the ingested logs, and applies whole graph embedding and sequence-based learning to capture rich structural properties of them. The key GNN technique leveraged by ProGrapher is graph2vec. ShadeWatcher (Zengy et al., 2022) is a recommendation-guided intrusion detector using provenance graphs. It borrows the recommendation concepts of user-item interactions into security concepts of system entity interactions and analyzes cyber threats in an automated and adaptive manner. threaTrace (Wang et al., 2022) emerges as an online approach dedicated to detecting host-based threats at the node level. Its GNN model is a tailored GraphSAGE (Hamilton et al., 2017) for learning rich contextual information in provenance graphs. MAGIC (Jia et al., 2024) leverages Graph Attention Network (GAT) (Velickovic et al., 2017) as its graph representation module. MAGIC employs masked graph representation learning to incorporate the capability of pre-training. It can adapt to concept drift with minimal computational overhead, making it applicable to real-world online APT detection. Flash (Rehman et al., 2024) is a comprehensive and scalable approach on data provenance graphs to overcome the limitations in accuracy, practicality, and scalability. Flash incorporates a novel adaptation of a GNN-based contextual encoder to encode both local and global graph structures into node embeddings efficiently. R-caid (Goyal et al., 2024) first incorporates root cause analysis into PIDS. Before training GNNs, R-caid links nodes to their root causes to build a new graph, intending to prevent it from mimicry and evasion attacks. Argus (Xu et al., 2024b) finds the performance of the prior IDS is questionable on large scale. It thus devises a form of discrete temporal graph and uses encoder-decoder unsupervised learning to detect different types of attacks. TAPAS (Zhang et al., 2025) leverages a stacked LSTM-GRU model and a task-guided segmentation algorithm to reduce the spatiotemporal dimensions of APT detection, achieving efficient, low-cost, and accurate detection. In addition to the aforementioned detectors, recent researchers have developed numerous useful tools for better understanding audit logs, such as data visualization analysis tool (Liu et al., 2025a) and counterfactual-driven attack explanation generator (Wu et al., 2025a).
@@ 5.2.2. Application Log-based Detectors
For traditional learning, Wei et al. (Xu et al., 2009) propose a general methodology to mine rich semantic information in console logs to detect large-scale system problems. Bodik et al. (Bodik et al., 2010) leverage a logistic regression model on a new and efficient representation of a datacenter’s state called fingerprint to detect previously seen performance crises in that datacenter. AMOD (Dong et al., 2018) uses the SVM HYBRID strategy to filter query annotations from web request logs and then update the stacked generalization detection model to efficiently detect web code injection attacks and obtain malicious queries to update the web application firewall (WAF) library.
@@
Due to the similarity between application logs and natural language texts, sequence neural networks such as Recurrent Neural Network (Hochreiter and Schmidhuber, 1997) and Transformer (Vaswani et al., 2017; Devlin et al., 2019) are widely employed. DeepLog (Du et al., 2017) employs LSTM to model system logs as natural language sequences. It is able to automatically learn benign log patterns and detect anomalies when there is a deviation between log patterns and the trained model. LogRobust (Zhang et al., 2019) finds previous methods do not work well under the close-world assumption and utilizes an attention-based LSTM model to handle unstable log events and sequences. LogAnomaly (Meng et al., 2019) identifies previous studies tend to cause false alarms by using indexes rather than semantics of log templates. Empowered by a novel, simple yet effective method termed template2vec, LogAnomaly is proven to successfully detect both sequential and quantitive log anomalies simultaneously. LogC (Yin et al., 2020) is a new log-based anomaly detection approach with component-aware analysis. It feeds both log template sequences and component sequences to train a combined LSTM model for detecting anomalous logs. NeuralLog (Le and Zhang, 2021) targets the performance caused by log parsing errors such as out-of-vocabulary words and semantic misunderstandings and employ BERT to perform neural representation. PLELog (Yang et al., 2021) is a semi-supervised anomaly detection approach that can get rid of time-consuming manual labeling and incorporate the knowledge on historical anomalies. SpikeLog (Qi et al., 2023) adopts a weakly supervised approach to train an anomaly score model, with the objective of handling a more reasonable premise scenario where a large number of logs are unlabeled. LogCraft (Zhang et al., 2024) is an end-to-end unsupervised log anomaly detection framework based on automated machine learning, which mitigates the cost of understanding datasets and makes multiple attempts for building algorithms. Tweezers (Cui et al., 2024) uses a large language model to identify entities and build a relationship graph, and generates embeddings through graph attention network optimization to achieve security incident detection. LogSer (Chai et al., 2024) parses logs by preprocessing parameters, splitting logs, tree parsing, and template merging. It then inputs relevant embeddings into BERT training to detect anomalies, generate reports, and perform incremental updates. LogDLR (Zhou et al., 2025) uses SBERT embeddings and a Transformer autoencoder with domain adversarial training to learn domain-invariant features, detecting anomalies via reconstruction error.
@@ 5.2.3. Network Traffic-based Detectors
Given the fact that network traffic is usually encrypted for secure communications, feature engineering-guided machine learning is widely applied in NIDS. NetPro (Li et al., 2017) employs traceability reasoning with Merkle Hash Trees and digital signatures to detect direct and indirect MANET routing attacks while preserving node privacy, and outputs a traceability graph to identify malicious nodes and behaviors. CATH (Guo et al., 2019) is a catastrophe-theory-based approach for DoS detection in software-defined networks (SDNs), which leverages the selection, normalization, and fusion of statistical flow attributes to model network states. Whisper (Fu et al., 2021) pays attention to both high accuracy and high throughput by utilizing frequency domain features. SigML++ (Trivedi et al., 2023) is an extension of SigML for supervised anomaly detection approach. SigML++ employs Fully Homomorphic Encryption and Artificial Neural Network (ANN) for detection, resulting in execution without decrypting the logs. OADSD (Zhang et al., 2023a) achieves task independently and has the ability of adapting to the environment over SD-WAN by using On-demand Evolving Isolation Forest. LtRFT (Tang et al., 2023) innovatively introduces Learning-To-Rank scheme for mitigating the low-rate DDoS attacks targeted at flow tables. AGC (Wu et al., 2025c) maps the original data into the embedding space through embedding learning to obtain more representative anchor points, thus achieving fine-grained classification of low-quality label data.
@@
In network traffic, packets consist of various contents and their flows can be represented as graphs. As a result, both graph neural network and sequence neural network are adopted in NIDS. Kitsune (Mirsky et al., 2018) is a plug and play NIDS that is allowed to detect attacks efficiently on the local network without supervision. It alleviates the problem that network gateways and router devices simply do not have the memory or processing power. MT-FlowFormer (Zhao et al., 2022a) is a semi-supervised framework to mitigate the lack of a mechanism for modeling correlations between flows and the requirement of a large volume of manually labeled data. I2RNN (Song et al., 2022) is an incremental and interpretable RNN for encrypted traffic classification, which can be efficiently adapted for incremental traffic types. ERNN (Zhao et al., 2022b) represents error-resilient RNN, which is a robust and end-to-end RNN model specially designed against network-induced phenomena. Euler (King and Huang, 2023) accelerates the most memory-intensive part, message-passing stage within GNN, with several concurrently-executed replicated GNNs. pVoxel (Fu et al., 2023) is an unsupervised method that proposes to leverage point cloud analysis to reduce false positives for the previous NIDS such as Whisper and Kitsune without requiring any prior knowledge on the alarms. NetVigil (Hsieh et al., 2024) is specially designed for east-west traffic within data center networks. It utilizes E-GraphSage and contrastive learning techniques to strengthen its resilience. Exosphere (Fu et al., 2024) detects flooding attacks by analyzing packet length patterns, without investigating any information in encrypted packets. DFNet (Zhao et al., 2024b) is a DDoS prevention paradigm denoted by preference-driven and in-network enforced shaping. RFH-HELAD (Zhong et al., 2024) consists of a $K$ classification model based on a deep neural network and a $K+1$ classification combining GAN and Deep kNN for detecting anomalies in network traffic. ReTrial (Zhao et al., 2024a) employs an improved graph attention network with Bayesian and EM algorithms to iteratively correct misleading links, enabling robust detection of encrypted malicious traffic. HEN (Wei et al., 2024) uses SMOTE to enhance data, trains LightGBM, generates explanations via SHAP, trains AE-LSTM to reconstruct SHAP values, sets a threshold from training errors, and marks test traffic with excess errors as attacks for intrusion detection. TCG-IDS (Wu et al., 2025b) is the first self-supervised temporal contrastive GNN for network intrusion detection, capturing spatiotemporal traffic dependencies with high accuracy and low false alarms. A-NIDS (Zha et al., 2025) uses a shallow fully connected network for real-time detection and a Stacked CTGAN generator to address catastrophic forgetting and old data storage costs. GTAE-IDS (Ghadermazi et al., 2025) uses a graph autoencoder with a Transformer encoder and DNN decoder to learn benign traffic, enabling label-free, near-real-time intrusion detection and new attack identification.